# Sleeping Kelly is a Thirder

### v1

@@ 6 Vulnerability of Halfers
Suppose Sleeping Beauty only accepts bets with wealth multiplier greater than 1 but she assigns a probability $p=\frac{1}{2}$ to the coin landing heads each time she is woken. We show that these “halfers” are vulnerable to Dutch books.
@@
Sleeping Beauty is vulnerable to Dutch books if she assigns a probability of $\frac{1}{2}$ to the coin landing heads or tails each time she is woken, even if she only accepts bets with a growth rate greater than 1.
@@
Now a Dutch book is possible. For example we can offer the “halfer” the following bets:
@@
7 Conclusion: Sleeping Kelly is a “thirder”
We approach the Sleeping Beauty problem by allowing Sleeping Beauty to place bets based on her beliefs. However, in contrast to canonical decision theory, we do not assume that Sleeping Beauty’s betting strategy is based on maximizing expected value. Instead, we assume that Sleeping Beauty aims to maximize the growth rate of her wealth, invoking the Kelly Criterion under multiplicative dynamics. This leads us to the conclusion that before the coin is flipped, Sleeping Beauty should bet based on a probability of $\frac{1}{2}$ that the coin lands tails, but each time Sleeping Beauty is woken she should bet based on a probability of $\frac{1}{3}$ that the coin lands tails.
Another approach to addressing the Sleeping Beauty problem is by examining whether Sleeping Beauty’s bets can be subjected to a Dutch book as a condition for rationality. We show that if Sleeping Kelly is a “thirder” and only accepts bets with a wealth multiplier greater than 1, she is not vulnerable to a Dutch book. However, “halfers” are vulnerable to Dutch books, even if they only accept bets they expect to have a wealth multiplier greater than 1.
Myopically maximizing the expected value of individual gambles is suboptimal for Sleeping Beauty, vulnerable to Dutch books, and arguably irrational.

### v2

@@ 8 Vulnerability of Halfers
If Sleeping Kelly is a halfer then she is vulnerable to Dutch books.
@@
Sleeping Beauty is vulnerable to Dutch books if she assigns a probability of $\frac{1}{2}$ to the coin landing heads or tails each time she is woken, even if she only accepts bets with a wealth multiplier greater than 1.
@@
Now a Dutch book is possible. For example we can offer the halfer the following bets for any suitably small $\epsilon$:
@@
9 Conclusion: Sleeping Kelly is a thirder
If Sleeping Beauty is allowed to choose bets on the outcome of the coin flip before being put to sleep and each time she is awakened, then she maximizes the growth rate of her wealth as a thirder sizing bets using the Kelly criterion under multiplicative wealth dynamics. This stands in contrast to the assumption that a bettor rationally maximizes the expected payoff of their bets. Furthermore, thirders cannot be subjected to diachronic Dutch books, while halfers remain vulnerable.